OpenStudy (nightmarenight):

How many solutions are there to the following system of equations? 6x + 12y = 11 4x + 8y = 3 0 1 2 infinitely many

OpenStudy (amoodarya):

multiply 1st by -2 multiply 2nd by +3 then add the equations to emit x and find y

OpenStudy (amoodarya):

6x + 12y = 11 --->-12x-24y=-22 4x + 8y = 3 ---->12x+24y=12 if add them 0=-22+12 so the system of equation has no solution